# EagleNet: Energy-Aware Fine-Grained Relationship Learning Network for Text-Video Retrieval

### 26 Mar 2026 v1

@@ Abstract
Text-video retrieval tasks have seen significant improvements due to the recent development of large-scale vision-language pre-trained models.
Traditional methods primarily focus on video representations or cross-modal alignment, while recent works shift toward enriching text expressiveness to better match the rich semantics in videos. However, these methods use only interactions between text and frames/video, and ignore rich interactions among the internal frames within a video, so the final expanded text cannot capture frame contextual information, leading to disparities between text and video. In response, we introduce Energy-Aware Fine-Grained Relationship Learning Network (EagleNet) to generate accurate and context-aware enriched text embeddings. Specifically, the proposed Fine-Grained Relationship Learning mechanism (FRL) first constructs a text-frame graph by the generated text candidates and frames, then learns relationships among texts and frames, which are finally used to aggregate text candidates into an enriched text embedding that incorporates frame contextual information. To further improve fine-grained relationship learning in FRL, we design Energy-Aware Matching (EAM) to model the energy of text-frame interactions and thus accurately capture the distribution of real text-video pairs. Moreover, for more effective cross-modal alignment and stable training, we replace the conventional softmax-based contrastive loss with the sigmoid loss. Extensive experiments have demonstrated the superiority of EagleNet across MSRVTT, DiDeMo, MSVD, and VATEX. Codes are available at https://github.com/draym28/EagleNet.
@@ 3.2.2 Energy-Aware Matching
Table 1: Performance on MSRVTT, boldface is the best results and underline is the runner-ups. Note that we have resolved the data-leakage problem in the original TMASS codes, and re-implemented XPool, CLIP-VIP, TMASS, TV-ProxyNet, and Video-ColBERT using the CLIP4Clip codebase to align with other methods. denotes the reproduced results, all of which are fine-tuned on ViT-B/32 or ViT-B/16.
@@
TV-ProxyNet [51] 46.8 72.1 81.3 46.6 76.0 85.0 58.9 89.1 94.4 650.2
Video-ColBERT [38] 40.2 68.6 79.3 43.9 73.9 83.5 57.1 87.4 93.1 627.0
@@
TV-ProxyNet [51] 47.9 75.2 84.5 49.7 79.7 88.0 64.0 91.7 95.9 676.6
Video-ColBERT [38] 44.5 71.7 82.1 47.7 78.3 86.6 61.1 90.2 95.3 657.5
@@ E.1 Datasets
(1) MSRVTT [52] includes 10,000 video clips, each of which has 20 captions. We use the 1k-A split following [31, 14] to use 9k videos for training and 1k videos for validation and test.
@@ E.4 Implementation Details
The pre-trained CLIP model (ViT-B/32 and ViT-B/16) is used as image and text encoder [37]. We use CLIP4Clip [33] as backbone, and apply similar configurations, such that the hidden dimension is $d=512$, weight decay $0.2$, dropout $0.3$, learning rate $10^{-7}$ for CLIP modules and $10^{-4}$ for non-CLIP modules; and we use a cosine scheduler for warmup, with the proportion of $0.1$. We uniformly sample $M=12$ frames for each video clips on all the datasets ($64$ for DiDeMo). All the frames are resized into $224\times 224$. We set the max length of text to $32$ for all the datasets ($64$ for DiDeMo). For FRL, the number of text candidates is $S=20$, the number of attention heads in RGAT is $H=4$, and the number of RGAT layers is $L=2$. For MCMC sampling in EAM, we set the sampling number to be the same as batch size $B$, step number $K=20$, step size $\eta=1$, and noise variance $\sigma^{2}=0.005$. The sigmoid loss is initialized as $\tau_{p}=4.77$ and b=12.93b=-12.93 following [56]. For loss coefficient, we set $\lambda_{\text{sup}}=0.8$ following [43], and $\lambda_{\text{eam}}=1.0$. In the training stage, we train all models on two A100 GPUs for $5$ epochs with a total batch size of $B=64$. In inference, we run on one single A100 GPU with a batch size of $8$.

### 01 Apr 2026 v3

@@ Abstract
Text-video retrieval tasks have seen significant improvements due to the recent development of large-scale vision-language pre-trained models. Traditional methods primarily focus on video representations or cross-modal alignment, while recent works shift toward enriching text expressiveness to better match the rich semantics in videos. However, these methods use only interactions between text and frames/video, and ignore rich interactions among the internal frames within a video, so the final expanded text cannot capture frame contextual information, leading to disparities between text and video. In response, we introduce Energy-Aware Fine-Grained Relationship Learning Network (EagleNet) to generate accurate and context-aware enriched text embeddings. Specifically, the proposed Fine-Grained Relationship Learning mechanism (FRL) first constructs a text-frame graph by the generated text candidates and frames, then learns relationships among texts and frames, which are finally used to aggregate text candidates into an enriched text embedding that incorporates frame contextual information. To further improve fine-grained relationship learning in FRL, we design Energy-Aware Matching (EAM) to model the energy of text-frame interactions and thus accurately capture the distribution of real text-video pairs. Moreover, for more effective cross-modal alignment and stable training, we replace the conventional softmax-based contrastive loss with the sigmoid loss. Extensive experiments have demonstrated the superiority of EagleNet across MSRVTT, DiDeMo, MSVD, and VATEX. Codes are available at https://github.com/draym28/EagleNet.
@@ 3.2.2 Energy-Aware Matching
Table 1: Performance on MSRVTT, boldface is the best results and underline is the runner-ups. Note that we have resolved the data-leakage problem in the original TMASS codes, and re-implemented XPool, CLIP-VIP, TMASS, TV-ProxyNet, and Video-ColBERT using CLIP4Clip codebase to align with other methods. denotes the reproduced results, all of which are fine-tuned on ViT-B/32 or ViT-B/16.
@@
TV-ProxyNet†∗ [51] 46.8 72.1 81.3 46.6 76.0 85.0 58.9 89.1 94.4 650.2
Video-ColBERT†∗ [38] 40.2 68.6 79.3 43.9 73.9 83.5 57.1 87.4 93.1 627.0
@@
TV-ProxyNet†∗ [51] 47.9 75.2 84.5 49.7 79.7 88.0 64.0 91.7 95.9 676.6
Video-ColBERT†∗ [38] 44.5 71.7 82.1 47.7 78.3 86.6 61.1 90.2 95.3 657.5
@@ E.1 Datasets
(1) MSRVTT [52] includes 10,000 video clips, each of which has 20 captions. We employ the 1k-A split [31, 14] to use 9k videos for training and 1k videos for validation and test.
@@ E.4 Implementation Details
The pre-trained CLIP model (ViT-B/32 and ViT-B/16) is used as image and text encoder [37]. We use CLIP4Clip [33] as backbone, and apply similar configurations, such that the hidden dimension $d=512$, weight decay $0.2$, dropout $0.3$, learning rate $10^{-7}$ for CLIP modules and $10^{-4}$ for non-CLIP modules; and we use a cosine scheduler for warmup, with the proportion of $0.1$. We uniformly sample $M=12$ frames for each video clip on all datasets ($64$ for DiDeMo). All frames are resized into $224\times 224$. We set the max length of text to $32$ for all datasets ($64$ for DiDeMo). For FRL, the number of text candidates $S=20$, the number of attention heads in RGAT $H=4$, and the number of RGAT layers $L=2$. For MCMC sampling in EAM, step number $K=20$, step size $\eta=1$, and noise variance $\sigma^{2}=0.005$. The sigmoid loss is initialized as $\tau_{p}=4.77$ and b=12.93b=-12.93 following [56]. For loss coefficient, we set $\lambda_{\text{sup}}=0.8$ following [43], and $\lambda_{\text{eam}}=1.0$ by default. We train all models on two A100 GPUs for $5$ epochs with a total batch size of $B=64$. In inference, we run on one A100 GPU with a batch size of $8$.